Add and simplify if possible.
9/(y-5) + 35/(y-5)^2 =

Answer by user_dude2008(1861)

9/(y-5) + 35/(y-5)^2 = 9(y-5)/(y-5)^2 + 35/(y-5)^2 = [9(y-5) + 35]/(y-5)^2 = (9y-45+35)/(y-5)^2 = (9y-10)/(y-5)^2